OpenStudy (anonymous):

Need the answer. MEDAL FOR REWARD The slope of any line perpendicular to , where A(0, -3) and B(4, 9) are two points on the line, is m = -3 -1/3 3

OpenStudy (anonymous):

m = (y2 - y1)/(x2-x1) where (x2,y2) = A and B = (x1, y1)

OpenStudy (anonymous):

IS THE ANSWER NUMBER 1 ?

OpenStudy (anonymous):

9-(-3)/4-0 = 3

OpenStudy (anonymous):

thanks